### v3

# Complements of caustics of the real $J_{10}$ singularities

Abstract.
The isotopy classification of Morse perturbations of parabolic real function singularities is completed.
1991 Mathematics Subject Classification:
Primary: 14P99. Secondary: 14Q30, 14B07
This work was supported by the Absorption Center in Science of the Ministry of Immigration and Absorption of the State of Israel
1. Introduction
$1.$
This work completes the isotopy classification of Morse perturbations of real parabolic function singularities. The analogous problem for simple singularities was solved in [10], [11], and [14]. For parabolic singularities of classes $X_{9}$ and $P_{8}$, this was done in [15] and [16], respectively. Now, we do the same for the remaining $J_{10}$ singularities. We prove that there are exactly 59 and 56 isotopy classes of Morse perturbations of $J_{10}^{1}$ and $J_{10}^{3}$ singularities, respectively.
In § 2, we describe a combinatorial invariant of isotopy classes of Morse functions that was used in previous works. Then, we compute all admissible values of this invariant for $J_{10}$ singularities and prove that each value can be realized by exactly one or two isotopy classes, depending on whether these classes are invariant or not under a certain symmetry. Finally, we investigate this symmetry condition. Additionally, we realize many isotopy classes by concrete polynomials.
The main invariant of isotopy classes is formulated in the terms of a graph, whose vertices are the collections of certain topological characteristics of particular Morse functions, and whose edges correspond to their standard surgeries. The invariant’s values are the subgraphs of this graph, into which it splits when the edges corresponding to surgeries changing the Morse isotopy class are removed. Our main criterion for the symmetry condition of isotopy classes can be expressed in the terms of one-dimensional cocycles of these subgraphs.
The most interesting part of the classification table obtained is the set of the isotopy classes consisting of polynomials with the maximal possible number of real critical points (ten), and only two different Morse indices. Similar to the $X_{9}$ and $P_{8}$ singularities, almost all of these classes are related to the splittings of the original function singularity into pairs of real critical points with the sum of Milnor numbers equal to ten. The only additional quadruple of isotopy classes (all of which are mapped to each other by certain symmetries) is related in the same way to the canonical extended Dynkin diagram of class $\tilde{E}_{8}$ (which is another name for $J_{10}$ singularities). This situation repeats that of cases $X_{9}$ and $P_{8}$, where the unique exceptional classes are related to the diagrams of types $\tilde{E}_{7}$ and $\tilde{E}_{6}$.
1.2. Main objects and definitions
A function singularity is a germ of a $C^{\infty}$-smooth function at a point where its differential vanishes. A smooth function $f:{\mathbb{R}}^{2}\to{\mathbb{R}}$ has a $J_{10}$ singularity at a point $a\in{\mathbb{R}}^{2}$ if in some local coordinates centered at this point it is quasihomogeneous of degree six with weights $\deg x=2$ and $\deg y=1$, and its Milnor number is finite (and then necessarily equal to ten, see [3], [4]). The level set $\{x:f(x)=f(a)\}$ of such a singularity in ${\mathbb{R}}^{2}$ can consist of one or three local branches at $a$. The corresponding subclasses are denoted by $J_{10}^{1}$ and $J_{10}^{3}$.
Table 1. Real simple and $J_{10}$ singularities in two variables
Notation Normal form Restriction
$A_{2k-1}$ $\pm x^{2k}\pm y^{2}$ $k\geq 1$
$A_{2k}$ $x^{2k+1}\pm y^{2}$ $k\geq 1$
$D_{2k}^{\pm}$ $x^{2}y\/\pm y^{2k-1}$ $k\geq 2$
$D_{2k-1}$ $\pm(x^{2}y\/+y^{2k-2})$ $k\geq 3$
$E_{6}$ $x^{3}\pm y^{4}$
$E_{7}$ $x^{3}+xy^{3}$
$E_{8}$ $x^{3}+y^{5}$
$J_{10}^{1}$ $\frac{1}{3}(x^{2}+y^{4})(x-\gamma y^{2})$ $\gamma\in(-\infty,\infty)$
$J_{10}^{3}$ $\frac{1}{3}(x^{2}-y^{4})(x-\gamma y^{2})$ $\gamma\in(-1,1)$
The normal forms of $J_{10}$ singularities, to which they can be reduced by a choice of local coordinates, are shown in the last two rows of Table 1. Other rows of this table show the normal forms of simple singularities into some of which the $J_{10}$ singularities can split by small perturbations.
The canonical versal deformations (see [3]) of function singularities of classes $J_{10}^{1}$ and $J_{10}^{3}$ consist respectively of polynomials
(1) $\frac{1}{3}(x-\gamma y^{2})(x^{2}+y^{4})+\lambda_{1}+\lambda_{2}y+\lambda_{3}y^{2}+\lambda_{4}y^{3}+\lambda_{5}y^{4}+\\ \lambda_{6}x+\lambda_{7}xy+\lambda_{8}xy^{2}+\lambda_{9}(x^{2}+3y^{4})y,\qquad\gamma\in(-\infty,+\infty)\ )$
and
(2) $\frac{1}{3}(x-\gamma y^{2})(x^{2}-y^{4})+\lambda_{1}+\lambda_{2}y+\lambda_{3}y^{2}+\lambda_{4}y^{3}+\lambda_{5}y^{4}+\\ \lambda_{6}x+\lambda_{7}xy+\lambda_{8}xy^{2}+\lambda_{9}(x^{2}-3y^{4})y,\qquad\gamma\in(-1,1),$
with ten parameters $\lambda_{1},\dots,\lambda_{9},$ and $\gamma$, see [7].
The set of non-Morse functions divides these families of polynomials into connected components called isotopy classes of Morse perturbations of $J_{10}$ singularities. Our main goal is the enumeration of these components.
As in [15], [16], we primarily consider slightly greater spaces than the canonical miniversal deformations (1) and (2), which can be reduced to these deformations by an appropriate group of coordinate changes.
Namely, we consider the 16-dimensional space of linear combinations of monomials $x^{\alpha}y^{\beta}$, such that $2\alpha+\beta\leq 6$, with a positive coefficient on the monomial $x^{3}$ and a non-degenerate principal quasihomogeneous part (consisting of monomials with $2\alpha+\beta=6$). This principal part can be of class $J_{10}^{1}$ or $J_{10}^{3}$. The spaces of polynomials with these principal parts are denoted, respectively, $\Phi_{1}$ and $\Phi_{3}$. Again, the sets of Morse functions of classes $\Phi_{1}$ and $\Phi_{3}$ split into connected components (= isotopy classes).
Definition 1.
The group $\mathfrak{G}$} consists of diffeomorphisms ${\mathbb{R}}^{2}\to{\mathbb{R}}^{2}$ of the form
(3) $\tilde{x}=ax+by+cy^{2}+\xi,\ \ \tilde{y}=dy+\eta,$
where $a>0,d>0.$ All coordinate systems in ${\mathbb{R}}^{2}$ obtained from the standard system by the group $\mathfrak{G}$} are called admissible.
The group $\mathfrak{G}$} acts freely on the spaces $\Phi_{1}$ and $\Phi_{3}$ and preserves the set of Morse functions. This group is homotopically trivial (i.e., is diffeomorphic to ${\mathbb{R}}^{6}$), and it reduces all functions of class $\Phi_{1}$ or $\Phi_{3}$ to normal forms (1) or (2).
Proposition 1.
Each orbit of the action of the group $\mathfrak{G}$} on the space $\Phi_{1}$ ((respectively, $\Phi_{3})$ intersects the space of all polynomials of type ((1)) ((respectively, ((2)))) transversally at a single point.
Proof. The transversality follows from the versality condition of our deformations. Let $f$ be a polynomial of the form (1), and let $\tilde{f}\equiv f\circ G$ for some diffeomorphism $G\in{\mathfrak{G}}$. If the coefficients $a$ and $d$ of this diffeomorphism (see (3)) are not equal to 1 or $c$ is not equal to 0, then the principal quasihomogeneous part of $\tilde{f}$ does not have the standard form of Table 1. If $a=d=1,$ $c=0$, but $\eta\neq 0$ then the coefficient of $\tilde{f}$ at the monomial $xy^{2}$ is non-trivial. If also $\eta=0$ but $b\neq 0$ then the coefficient at $y^{5}$ is not equal to thrice the coefficient at $x^{2}y$. If $\eta=0$ but $\xi\neq 0$, then the coefficient at $x^{2}$ is non-trivial. Thus, $\tilde{f}$ belongs to deformation (1) only if $G=\mbox{Id}$. The proof for $J_{10}^{3}$ is analogous. $\Box$
So, each space $\Phi_{1}$ or $\Phi_{3}$ is canonically diffeomorphic to the direct product of the group $\mathfrak{G}$} and the space of polynomials (1) or (2) ((which is diffeomorphic to ${\mathbb{R}}^{10})$. Additionally, the isotopy classes of Morse functions in spaces $\Phi_{1}$ and $\Phi_{3}$ are in a one-to-one correspondence with the connected components of their intersections with the spaces of polynomials (1) and (2), respectively.
The group $\mathfrak{G}$} also reduces any function of class $\Phi_{1}$ or $\Phi_{3}$ to the normal form with zero coefficients at the monomials $x^{2},x^{2}y,$ and $x^{2}y^{2}$, coefficient 1 at $x^{3}$, and either coefficient $\pm 1$ at $y^{6}$ and coefficient zero at $y^{5}$, or coefficient zero at $y^{6}$ and $xy^{3}$ and coefficient $\pm 1$ at $xy^{4}$.
The main result of the paper is a list of all isotopy classes of Morse polynomials of classes $\Phi_{1}$ and $\Phi_{3}$ or, equivalently, of the form (1) and (2). It is formulated in Theorems 13, 14, and 15.
As a byproduct, we list all possible splittings of real $J_{10}$ singularities into pairs of singular points, such that the sum of their Milnor numbers is ten.
Definition 2 (cf. [7], §2).
For any pair of simple singularity classes $\Xi$ and $\tilde{\Xi}$ with $\mu(\Xi)+\mu(\tilde{\Xi})=10$, the notation $\{\Xi+\tilde{\Xi}\}\rightsquigarrow J_{10}^{1}$ or $\{\Xi+\tilde{\Xi}\}\rightsquigarrow J_{10}^{3}$ means that there exists a smooth function $f:{\mathbb{R}}^{2}\to{\mathbb{R}}$ with a critical point of class $J_{10}^{1}$ or $J_{10}^{3}$ and a one-parametric deformation $F:{\mathbb{R}}^{2}\times[0,\varepsilon)\to{\mathbb{R}}$, $F(\cdot,0)\equiv f,$ of this function, such that, for any $\tau\in(0,\varepsilon)$, the corresponding function $f_{\tau}\equiv F(\cdot,\tau)$ has a critical point of class $\Xi$ and a critical point of class $\tilde{\Xi}$ in such a way that these two critical points depend continuously on $\tau$ and tend to the $J_{10}^{\ast}$ singular point of the function $f$ as $\tau$ tends to 0.
Table 2. Approximation of $J_{10}$ singularities by bisingularities
Type $J_{10}^{1}$ $J_{10}^{3}$
$E_{8}+A_{2}$ Yes No
$D_{8}^{+}+A_{2}$ No No
$D_{8}^{-}+A_{2}$ No Yes
$A_{8}+A_{2}$ Yes No
$E_{7}+A_{3}$ No Yes
$D_{7}+A_{3}$ Yes No
$A_{7}+A_{3}$ Yes Yes
Type $J_{10}^{1}$ $J_{10}^{3}$
$E_{6}+D_{4}^{+}$ Yes No
$E_{6}+D_{4}^{-}$ No No
$E_{6}+A_{4}$ Yes No
$D_{6}^{\pm}+D_{4}^{\pm}$ No No
$D_{6}^{+}+A_{4}$ Yes No
$D_{6}^{-}+A_{4}$ No Yes
Type $J_{10}^{1}$ $J_{10}^{3}$
$A_{6}+D_{4}^{+}$ No No
$A_{6}+D_{4}^{-}$ No Yes
$A_{6}+A_{4}$ Yes No
$A_{5}+A_{5}$ Yes Yes
$A_{5}+D_{5}$ Yes Yes
$D_{5}+D_{5}$ No Yes
Theorem 2 (see § 7).
For any pair of simple real singularity classes $\Xi$ and $\tilde{\Xi}$ with $\mu(\Xi)+\mu(\tilde{\Xi})=10$, we have $\{\Xi+\tilde{\Xi}\}\rightsquigarrow J_{10}^{1}$ ((respectively, $\{\Xi+\tilde{\Xi}\}\rightsquigarrow J_{10}^{3})$ if and only if “Yes” is written in Table 2 at the intersection of the row $\Xi+\tilde{\Xi}$ and the column $J_{10}^{1}$ ((respectively, $J_{10}^{3})$.
This theorem is a “real” analog of the corresponding part of Table III of [7].
2. Invariants of isotopy classes of Morse polynomials
This section essentially repeats Section 2 of [15].
2.1. Trivial invariant
The simplest invariant of isotopy classes of Morse polynomials in ${\mathbb{R}}^{2}$ is their passport, i.e. the triple $(m_{-},m_{\times},m_{+})$ of the numbers of their minima, saddles, and maxima.
By the index considerations, for any Morse polynomial of class $\Phi_{1}$ or $\Phi_{3}$ the sum $M\equiv m_{-}+m_{\times}+m_{+}$ is an even number no greater than ten, and the Euler number $m_{-}-m_{\times}+m_{+}$ is equal to 0 for all Morse polynomials of class $\Phi_{1}$ and to 2-2 for $\Phi_{3}$.
Therefore, when studying a particular class $\Phi_{1}$ or $\Phi_{3}$, we will express the passports by only pairs of numbers, $m_{+}$ and the total number $M$ of real critical points (which determine the remaining passport numbers).
2.2. Set-valued invariant and virtual Morse functions
Definition 3.
A polynomial $f:({\mathbb{C}}^{2},{\mathbb{R}}^{2})\to({\mathbb{C}},{\mathbb{R}})$ of class $\Phi_{1}$ or $\Phi_{3}$ is generic if it has only Morse critical points in ${\mathbb{C}}^{2}$, and all the corresponding critical values are different and not equal to 0.
We associate a set of discrete topological characteristics, called a virtual Morse function, with any generic Morse polynomial $f$.
12345678910
12345687109
Figure 1. Standard systems of paths
If $f$ is a generic polynomial of class $\Phi_{1}$ or $\Phi_{3}$, then the set $V_{f}\subset{\mathbb{C}}^{3}$ defined by the equation $f(x,y)+z^{2}=0$ is a smooth complex surface that is homotopy equivalent to the wedge of ten two-dimensional spheres (see, for example, [9] and [3]). The homology group $H_{2}(V_{f})$ is generated by vanishing cycles (see [3], [12]) that are defined by a system of non-intersecting paths in ${\mathbb{C}}^{1}$ connecting the non-critical value 0 with all critical values of $f$, see Fig. 1. We choose these paths so that those going to complex conjugate non-real critical values are symmetric about the real axis, and those going to real values lie in the upper half-plane where the imaginary parts are smaller than the absolute values of the imaginary parts of all non-real critical values.
Let us fix an orientation of ${\mathbb{R}}^{3}$ somehow. Then there is a canonical choice of orientations of all vanishing cycles defined by such a system of paths, see §V.1.6 of [12]. In particular, complex conjugation in ${\mathbb{C}}^{3}$ must take the oriented vanishing cycles defined by complex conjugate paths into each other with the coefficient 1 and not 1-1.
An order of these vanishing cycles can also be canonically defined. Specifically, the cycles that vanish at real critical points are listed in ascending order of the corresponding critical values.
Definition 4 (see [4], §V.3).
A virtual Morse function associated with a generic Morse polynomial $f:({\mathbb{C}}^{2},{\mathbb{R}}^{2})\to({\mathbb{C}},{\mathbb{R}})$ of class $\Phi_{1}$ or $\Phi_{3}$ is a collection of its topological data consisting of
a) the $10\times 10$ matrix of intersection indices in $V_{f}$ of canonically ordered and oriented vanishing cycles $\Delta_{i}\in H_{2}(V_{f})$ corresponding to all critical values of $f$ and defined by a system of paths as above,
b) the string of ten intersection indices in $V_{f}$ of these vanishing cycles with the naturally oriented set of real points, $V_{f}\cap{\mathbb{R}}^{3}$;
c) the string of positive Morse indices of all real critical points of the function $f(x,y)+z^{2}$, and
d) the numbers of negative and positive real critical values of $f$.
Example. Some four virtual Morse functions with eight real critical points are shown in Fig. 2. Two vertical lines in each table indicate the last element of the corresponding virtual Morse function: in all four cases, the number of negative, positive and non-real critical values are 3, 5, and 2, respectively. The real critical points of these functions are only minima and saddles.
[…]
Figure 2. Virtual Morse functions
Definition 5.
A critical point of a virtual Morse function is any column of its data set as in Fig. 2, i.e. a column of the intersection matrix, the intersection index with the set of real points, and a Morse index or the information that the critical point is non-real.
Remark 3.
If there are more than one of pairs of non-real critical values of $f$, then there can be more than one virtual Morse function associated with the same real polynomial, because the choice of a system of paths is not homotopically unique: see Fig. 1.
Definition 6.
Elementary virtual surgeries of virtual Morse functions include six transformations of their data, modeling the standard local topological surgeries of the corresponding real Morse polynomials, namely
  • $s1,s2$:
    collision of two neighboring real critical values at a non-zero value, after which the corresponding two critical points either ($s1$) meet and leave the real domain, or ($s2$) change the order in ${\mathbb{R}}^{1}$ of their critical values;
  • $s3,s4$:
    collision of two complex conjugate critical values at a point on the line ${\mathbb{R}}^{1}\setminus\{0\}$, after which the corresponding critical points either ($s3$) meet at a real point and enter real space, or ($s4$) miss each other in the complex domain, while the imaginary parts of their critical values change their signs;
  • $s5,s6$:
    jumps of real critical values up ($s5$) or down ($s6$) through 0;
                 and additionally
  • $s7$:
    specifically virtual transformations within the classes of virtual Morse functions associated with the same real Morse polynomials, that are caused by flips of standard systems of paths going from 0 to non-real critical values (see Fig. 1 and [3], and also Figs. 19–21 of volume 2).
The results of all these virtual surgeries are determined by the data of the initial virtual Morse functions. For a detailed description of these standard flips of data, see §V.8 of [12], the explicit formulas for them are given in the comments of the computer program that performs them, see the link on page 3.2. In particular, attempting to perform the surgery $s1$ or $s2$ over real critical values $v_{i}$ and $v_{i+1}$ begins with examining the intersection index $\langle\Delta_{i},\Delta_{i+1}\rangle$ of the corresponding vanishing cycles. If this index is 0, then surgery $s2$ occurs; if the index is 1, then surgery $s1$ occurs; in all other cases the surgery fails. Similarly, a collision of two complex conjugate critical values at a real point not separated from 0 by other critical values follows scenario $s4$ if the intersection index is 0, scenario $s3$ if the index is $1$ or 1-1, and fails in all other cases; in the second case the sign of the intersection index allows us to predict the Morse indices of the newborn real critical points.
We will denote by $s1,\dots,s6$ both the real surgeries of real Morse functions and the corresponding elementary virtual surgeries.
Remark 4.
If our deformed singularity is neither parabolic nor simple, then we cannot be sure that the real surgery corresponding to a virtual one can be realized any time when these intersection index conditions are satisfied. For simple singularities this realization is guaranteed by the properness of the so-called Lyashko–Looijenga map (see [8]). For parabolic singularities, it follows from the Jaworski’s work [6], [7], see Proposition 7 below.
Let $f:{\mathbb{R}}^{2}\to{\mathbb{R}}$ be a generic Morse polynomial of class $\Phi_{1}$ or $\Phi_{3}$.
Definition 7.
An (abstract) virtual Morse function of type $f$ is any collection of data as in Definition 4 (i.e. a matrix, two strings and two numbers) obtained from an arbitrary virtual Morse function associated with $f$ by an arbitrary chain of elementary virtual surgeries.
The formal graph of type $f$ is the graph, whose vertices are all virtual Morse functions of type $f$, and two such vertices are connected by an edge if and only if they can be obtained from each other by an elementary virtual surgery.
The virtual component $S(f)$ of the formal graph of type $f$ is its subgraph, whose vertices are only the virtual Morse functions of type $f$ that can be obtained from virtual Morse functions associated with $f$ via arbitrary chains of virtual surgeries $s2,s4,s5,s6$ and $s7$ from Definition 6 (i.e. all surgeries that do not model the collision of critical points).
Example 5.
Denote the four virtual Morse functions in Fig. 2 by the letters A, B, C, and D, as shown in the diagram $\left|{A\ B\atop C\ D}\right|$. The transition $A\to B$ describes the passage $s4$ of two imaginary critical values between the fifth and sixth real critical values. The transitions $A\to C$, $B\to D$ and $C\to D$ are respectively the surgeries $s2$ at which the second and third, the seventh and eighth, and the fourth and fifth real critical values meet and overtake each other. These four elementary surgeries form a commutative diagram, and the chain of passages $A\to B\to D\to C\to A$ forms a cycle in a virtual component of the formal graph.
Proposition 6.
For all generic polynomials $f$ of the same class $\Phi_{1}$ or $\Phi_{3}$ the formal graphs of type $f$ are the same.
If two generic Morse polynomials $f$ and $\tilde{f}$ belong to the same connected component of the set of Morse polynomials of class $\Phi_{1}$ or $\Phi_{3}$, then their virtual components, $S(f)$, and S(f~)S($\tilde${f}), are the same.
Proof. The first statement follows from the connectedness of the spaces $\Phi_{1}$ and $\Phi_{3}$: any two real Morse polynomials from such a space can be connected by a path that intersects the variety of non-Morse polynomials at finitely many points, each of which corresponds to standard surgery. The second statement follows immediately from the definitions. $\Box$
Proposition 6 justifies the following definitions.
Definition 8.
The formal graph of a (not necessarily Morse) polynomial ${\mathbb{R}}^{2}\to{\mathbb{R}}$ of class $\Phi_{1}$ or $\Phi_{3}$ is the formal graph of an arbitrary generic Morse polynomial of the same class.
The virtual component of a (not necessarily strictly) Morse polynomial of class $\Phi_{1}$ or $\Phi_{3}$ is the virtual component of an arbitrary generic polynomial from the same connected component of the space of all Morse polynomials of this class.
Proposition 7.
For any generic Morse polynomial $f$ of class $\Phi_{1}$ or $\Phi_{3}$, a virtual Morse function $\varphi$ associated with it, and a virtual Morse function $\tilde{\varphi}$ connected with $\varphi$ by an edge of the formal graph, there exists a generic Morse polynomial $\tilde{f}$ and a path in the space $\Phi_{1}$ or $\Phi_{3}$ connecting $f$ and $\tilde{f}$ and containing only one non-generic point at which it experiences a standard surgery of the same type as the edge $\varphi,\tilde{\varphi}$].
Proof. In the case of class $\Phi_{3}$, the proof repeats the proof of Proposition 2 in [15] for $X_{9}^{+}$ singularities. The only difference is that it considers the canonical complex versal deformation of $J_{10}$ singularities (having exactly the form (2)) instead of the canonical versal deformation of $X_{9}$. In the case $\Phi_{1}$, we reduce our study to this complex deformation using the substitution $x=e^{\pi i/4}\tilde{x}$. $\Box$
Corollary 8.
Let $f$ be a generic polynomial of class $\Phi_{1}$ or $\Phi_{3}$, then
a)) every virtual Morse function of type $f$ is associated with a generic real polynomial $\tilde{f}$ of the same class $\Phi_{1}$ or $\Phi_{3}$ as $f$;
b)) every virtual Morse function from the virtual component $S(f)$ is associated with some generic real polynomial from the same connected component as $f$ of the space of Morse polynomials of class $\Phi_{1}$ or $\Phi_{3}$;
Definition 9.
The set-valued invariant of a real Morse polynomial $f$ of class $\Phi_{1}$ or $\Phi_{3}$ is the set of virtual Morse functions corresponding to all vertices of the virtual component of $S(f)$.
The Card invariant of such a Morse polynomial $f$ is the cardinality of this set of vertices.
Clearly, the set-valued invariant determines the “passport” invariant (which can be read from the bottom line of any virtual Morse function representing it).
Proposition 9.
If two Morse polynomials of class $\Phi_{1}$ or $\Phi_{3}$ are associated with the same virtual Morse function, then they either belong to the same connected component of the space of generic polynomials of this class, or they belong to components that are mapped to each other by the involution
(4) $f(x,y)\leftrightarrow f(x,-y).$
The proof of this proposition will be given in § 4.
Corollary 10.
Each virtual component of the formal graph of type $\Phi_{1}$ or $\Phi_{3}$ is associated with one or two isotopy components of the space of Morse polynomials of the corresponding class $\Phi_{1}$ or $\Phi_{3}$. $\Box$
Definition 10.
An isotopy class of Morse polynomials and the associated virtual component of the formal graph of class $\Phi_{1}$ or $\Phi_{3}$ are called achiral (respectively, chiral) if the involution (4) takes this isotopy class to itself (respectively, to a different component).
Finally, our system of isotopy invariants of Morse polynomials of classes $\Phi_{1}$ and $\Phi_{3}$ consists of the set-valued invariant of Definition 9 and the reflection class in the case of chiral classes. By Proposition 9, this system of invariants separates all isotopy classes.
2.3. D-graph invariant
In this subsection we only consider the real polynomials $({\mathbb{C}}^{2},{\mathbb{R}}^{2})\to({\mathbb{C}},{\mathbb{R}}),$ whose critical points are all real. In this case the set-valued invariant of §2.2 has the following transparent interpretation.
Let $f$ be a generic polynomial ${\mathbb{R}}^{2}\to{\mathbb{R}}$ of class $\Phi_{1}$ or $\Phi_{3}$ with only real critical points, in particular all ten of its critical values are real and distinct, and 0 is a non-critical value of $f$. The matrix of intersection indices of vanishing cycles $\Delta_{i}\in H_{2}(V_{f})$ (numbered in the ascending order of the corresponding critical values) can be depicted by its Coxeter-Dynkin graph (see e.g. [3]) with ten ordered vertices corresponding to all critical values of $f$. Namely, if the intersection index $\langle\Delta_{i},\Delta_{j}\rangle$ is positive, then the corresponding vertices $v_{i}$ and $v_{j}$ are connected by $\langle\Delta_{i},\Delta_{j}\rangle$ solid segments; if $\langle\Delta_{i},\Delta_{j}\rangle$ is negative then they are connected by $\langle\Delta_{i},\Delta_{j}\rangle$ dashed segments.
Definition 11 (see [15]).
The D-graph of a generic real Morse polynomial $f$ with only real critical points is (the isomorphism class of) the oriented graph with vertices labeled with indices 0, 1 and 2, which is obtained from the Coxeter-Dynkin graph of $f$ by
1) orienting each edge of the graph from the vertex corresponding to the critical point with the lower critical value to the vertex corresponding to the critical point with the higher critical value;
2) labeling each vertex of the graph with the Morse index of the corresponding critical point of $f$, and
3) forgetting the ordering of the vertices.
Notation. In the pictures of the D-graphs (see Figs. 317), instead of numerical Morse indices, we will label the vertices corresponding to minima, saddlepoints and maxima by white circles, black circles, and white squares, respectively.
Remark 11.
The D-graph of a real Morse polynomial is determined by an arbitrary virtual Morse function associated with this polynomial. Indeed, the intersection matrix and the Morse indices are the elements of the virtual Morse function, and the orientation of the edges follows from the order of the rows and columns of this matrix determined by the order of the corresponding critical values. In this way also D-graphs of arbitrary abstract virtual Morse functions with only real critical points are well defined.
Theorem 12.
In the restriction to generic Morse polynomials with only real critical points, the D-graphs form an invariant of isotopy classes of Morse functions. This invariant is equivalent to the set-valued invariant from §2.2.
Proof of this theorem repeats the proof of Theorem 2 of [15].
2.4. The up-down involution
The involution
(5) $f(x,y)\leftrightarrow-f(-x,y)\$
acts on the spaces $\Phi_{1}$ and $\Phi_{3}$.
This action can also be extended to the corresponding virtual graphs. Namely, for any system of paths for the function $f(x,y)$ (see Fig. 1) we take the system of paths for f(x,y)-f(-x,y) obtained from it by the composition of the multiplication by 1-1 and the complex conjugation in ${\mathbb{C}}^{1}$. The virtual Morse function associated with the function f(x,y)-f(-x,y) and defined by the resulting system of paths is determined by that of the original virtual Morse function associated with $f(x,y)$, see § 2.2 in [16].
In particular, if two functions with ten real critical points are related by this involution, then their $D$-graphs are obtained from each other by replacing all minima by maxima and vice versa, and reversing the directions of all edges, cf. Proposition 16 in [16]
3. Enumeration of virtual components
Table 3. Numbers of virtual components of type $J_{10}^{1}$ (left) and $J_{10}^{3}$ (right)
[…]
0 1 1 1 1 1(2) 9(18)
1 0 1 1 1 1 1(2)
2 0 0 1 1 1 1
3 0 0 0 1 1 1
4 0 0 0 0 1(2) 1(2)
5 0 0 0 0 0 9(18)
$\Sigma$ 1 2 3 4 5(7) 22(42)
[…]
0 0 1 1 1 2 9(18)
1 0 0 1 1 1 2(3)
2 0 0 0 1 1 1(2)
3 0 0 0 0 2 2(3)
4 0 0 0 0 0 9(18)
$\Sigma$ 0 1 2 3 6 23 (44)
3.1.
Table 3 shows the number of virtual components and the number of isotopy classes of Morse polynomials of type $\Phi_{1}$ and $\Phi_{3}$ with any value of the passport $(M,m_{+})$. The second of these numbers is given in parentheses when it differs from the first one. This table is detailed in Theorems 13, 14, 15.
$\diamond$
Figure 3. $\Phi_{1}$, one maximum (122298)
$\diamond\diamond$
Figure 4. $\Phi_{1}$, two maxima (26378)
Theorem 13 ($\Phi_{1}$).
1. There are exactly 22 virtual components of type $\Phi_{1}$ consisting of virtual Morse functions with ten real critical points:
  • one component consisting of virtual Morse functions with exactly one local maximum, D-graph shown in Fig. 3 ((right)) and Card invariant equal to 122298,
  • one component of virtual Morse functions with exactly two local maxima, D-graph shown in Fig. 4 ((right)) and Card=26378,
  • nine components of virtual Morse functions without local maxima; their D-graphs are shown in Figs. 59, and their Card invariants are indicated in the captions of these figures,
  • eleven virtual components obtained from the ones listed above by the involution ((5)) of the space $\Phi_{1}$.
2. There are exactly five virtual components of type $\Phi_{1}$ with eight real critical points:
  • one component with two local maxima and two minima, and Card=10890;
  • one component with exactly one local maximum, and Card=27378,
  • one component without local maxima, having Card=166554,
  • two virtual components obtained from the last two by the involution ((5)).
3. There are exactly four virtual components of type $\Phi_{1}$ with six real critical points:
  • one component with exactly one local maximum and Card=10122,
  • one component without local maxima, having Card=44128,
  • two virtual components obtained from the previous two by the involution ((5)).
4. There are exactly three virtual components of type $\Phi_{1}$ with four real critical points:
  • one component with exactly one local maximum and Card=7850,
  • one component with no local maxima and Card=15850,
  • a virtual component obtained from the previous one by the involution ((5)).
5. There are exactly two virtual components of type $\Phi_{1}$ with two real critical points: one component without local maxima and one component without minima obtained from it by ((5)), both having Card=10608.
6. There is only one virtual component of type $\Phi_{1}$ without real critical points, its Card invariant is equal to 17642.
Figure 5. 97702 (left) and 93489 (right)
Figure 6. 68145 (left) and 42372 (right)
Figure 7. 52415 (left) and 63085 (right)
Figure 8. 82500 (left) and 27940 (right)
Figure 9. 33528
$\diamond$
Figure 10. $\Phi_{3}$, one maximum (77374)
$\diamond$
Figure 11. $\Phi_{3}$, one maximum (225148)
$\diamond\diamond$
Figure 12. $\Phi_{3}$, two maxima (128634)
Theorem 14 ($\Phi_{3}$).
1. There are exactly 23 virtual components of type $\Phi_{3}$ with ten real critical points:
  • two components with exactly one local maximum: one with Card=77374 and D-graph shown in Fig. 10 ((right)), and the other with Card=225148 and D-graph shown in Fig. 11 ((right));
  • one component with two local maxima, two minima, Card=128634 and D-graph shown in Fig. 12 ((right));
  • nine virtual components without local maxima, whose D-graphs are shown in Figs. 1317 and Card invariants are indicated in the captions of these pictures;
  • eleven virtual components obtained by the involution ((5)) from the components listed above ((except for the component with two minima and two maxima)).
2. There are exactly six virtual components of type $\Phi_{3}$ with eight real critical points:
  • one component with one local maximum, two minima and Card=66906,
  • two components with no local maxima and values of Card invariant equal to 131148 and 82350,
  • three virtual components obtained from these mentioned in the previous two items by involution ((5)).
3. There are exactly three virtual components of type $\Phi_{3}$ with six real critical points:
  • one component with one local maximum, one minimum and Card=26922,
  • one component without local maxima having Card=57442,
  • a virtual component obtained from the previous one by the involution ((5)).
4. There are exactly two virtual components of type $\Phi_{3}$ with four real critical points: one component without local maxima, having Card=21410, and one component obtained from the previous one by the involution ((5));
5. There is only one virtual component of type $\Phi_{3}$ with two real critical points: it has Card=14778 and no local extrema;
6. There are no virtual components of type $\Phi_{3}$ without real critical points.
Figure 13. 230472 (left) and 102234 (right)
Figure 14. 89320 (left) and 75108 (right)
Figure 15. 63756 (left) and 59862 (right)
Figure 16. 53130 (left) and 34870 (right)
Figure 17. 29370
Theorem 15.
1. All virtual components of the type $\Phi_{1}$ with ten real critical points are chiral except for those with exactly two maxima or two minima ((see Fig. 4)), which are achiral. All virtual components of the type $\Phi_{1}$ with less than ten real critical points are achiral except for the ones with eight real critical points, having no minima or no maxima.
2. All virtual components of the type $\Phi_{3}$ with ten real critical points are chiral except for those with exactly one maximum or one minimum ((see Fig. 10)), which are achiral. All virtual components of the type $\Phi_{3}$ with less than ten real critical points are achiral.
3.2. On the proofs of enumeration theorems
All possible values of the Card and D-graph invariants mentioned in Theorems 13 and 14 were found using the computer program described in [4], [13]. For any class $\Phi_{1}$ or $\Phi_{3}$, we use the Gusein-Zade–A’Campo method (see [1], [5]) to compute the intersection matrix of vanishing cycles of a Morse polynomial of this class having only real critical points. Then, using theorem 1.4 of §5.1 in [12], we calculate the intersection indices of these vanishing cycles with the set of real points. Thus, we obtain a virtual Morse function associated with this polynomial. Starting from this initial data, the combinatorial program
https://drive.google.com/file/d/19FK3NDqHr01CVyVegO6wNUaedrvi_8nJ/view?usp=sharing
runs through the entire formal graph of this type and counts, in particular, the number of all virtual Morse functions of this type and the number of virtual functions with each passport invariant value. For each passport value for which this number is not zero, it then (upon request) finds a virtual Morse function with this value. Starting from this virtual Morse function, a slightly modified version of the same program (with virtual surgeries of types $s1$ and $s3$ disabled) counts the number of virtual Morse functions in its virtual component, i.e. the Card invariant of this virtual function. For example, this program
https://drive.google.com/file/d/1bGV16NMqm-VxvChKxjLTqxORUtKeIVQe/view?usp=sharing
with certain initial data handles the virtual component of the $\Phi_{3}$ class, which has eight real critical points and Card = 131148. If this number is smaller than the total number of virtual functions with that passport, we find another virtual function with the same passport but in a different virtual component. Then we calculate its Card value as well. We continue this process until the sum of the different values of the Card invariant of virtual functions with any given passport reaches the total number of virtual functions with that passport. This computation (and the subsequent reconstruction of D-graphs from the virtual Morse functions representing these components) proves all statements of Theorems 13 and 14.
In sections 5 and 6, we investigate the chirality of all these components and prove Theorem 15. Specifically, in § 5 we realize nearly all achiral virtual components by polynomials that are symmetric with respect to the reflection (4). In § 6.1 we describe a homological criterion of chirality in the terms of the formal graph and use it to demonstrate the achirality of the remaining achiral components. In § 6.2, we prove the chirality of all chiral components.
Remark 16.
The spaces $\Phi_{1}$ and $\Phi_{3}$ are invariant under the involution (5). This involution maps Morse polynomials with any value $(M,m_{+})$ of the “passport” invariant to polynomials with the value $(M,M/2-m_{+})$ in the case $\Phi_{1}$ and to polynomials with the value $(M,M/2-1-m_{+})$ in the case $\Phi_{3}$. Therefore, it is sufficient to study only the virtual Morse functions with $m_{+}\leq M/4$ (respectively, $m_{+}\leq M/4-1$).
3.3. Normalization of $D$-graphs
Definition 12.
An edge of a $D$-graph is called normal if
a) it is directed from a vertex with a smaller Morse index to a vertex with a larger index,
b) it is dashed if the parities of Morse indices of corresponding critical points are the same, and is solid if these parities are different.
Otherwise, this edge is called a tunnel edge. The normalization of a $D$-graph consists in removing all its tunnel edges.
By analyzing Figs. 317, we see that
a) all $D$-graphs of polynomials with both maxima and minima are already normal,
b) seventeen out of eighteen $D$-graphs of polynomials without maxima shown in these figures split into pairs of standard Coxeter–Dynkin graphs of some simple singularities after normalization;
c) the remaining $D$-graph (see Fig. 13 left) splits into an isolated vertex and the extended Coxeter–Dynkin graph of type $\tilde{E}_{8}$.
A very similar situation occurs for $D$-graphs of other parabolic singularities, see [15] and [16] (with extended graphs of types $\tilde{E}_{7}$ and $\tilde{E}_{6}$).
4. Proof of Proposition 9
Lemma 17.
The subset of the space $\Phi_{1}$ consisting of polynomials having a singularity of class $E_{8}$ with zero critical value, and also a singularity of class $A_{2}$, is the union of four connected components diffeomorphic to ${\mathbb{R}}^{7}$ and mapped to each other by the involutions ((4)) and ((5)). Two of these components consist of polynomials with negative critical values at the $A_{2}$ points, and the other two consist of polynomials with positive values. Each component is swept out by a one-parametric family of orbits of the group $\mathfrak{G}$}. The intersection of each such component with the parameter space of the deformation ((1)) is an orbit of the group $\mathbb{R}$$_{+}$ of positive quasihomogeneous dilations
(6) $T_{t}:f(x,y)\to t^{-6}f(t^{2}x,ty),\quad t>0.$
Proof. Let $f\in\Phi_{1}$ be a polynomial with a critical point of class $E_{8}$, in particular its 2-jet at this point is trivial and 3-jet has the form $q(x-\nu y)^{3},$ $q>0$. The substitution $\tilde{x}=x-\nu y$ (which belongs to the group $\mathfrak{G}$}) reduces it to the form $\tilde{x}^{3}$. The resulting function has zero coefficients at the monomials $y^{4}$ and $\tilde{x}y^{3}$, because otherwise its critical point 0 would be of class $E_{6}$ or $E_{7}$. Conversely, the coefficient at the monomial $y^{5}$ is non-zero: otherwise the Milnor number is greater than $8$. So, this function is a linear combination of monomials $x^{3},$ $x^{2}y^{2}$, $xy^{4}$, $y^{6}$, and $y^{5}$. An additional substitution of the form $\check{x}=\tilde{x}-\eta y^{2}$ (which also belongs to the group $\mathfrak{G}$}) reduces $f$ to the form
(7) $\check{x}^{3}+a\check{x}y^{4}\pm y^{6}+\tau y^{5}\quad\mbox{or}\quad\check{x}^{3}\pm\check{x}y^{4}+\tau y^{5},\qquad\tau\neq 0.$
The condition of having also a critical point of class $A_{2}$ (that is, the coincidence of two non-zero solutions of the system $f^{\prime}_{x}=0=f^{\prime}_{y}$) prohibits the second possibility and implies the condition $a=0$ in the first possibility of the alternative (7). The resulting set of polynomials of the form
(8) $x^{3}\pm y^{6}+\tau y^{5}$
splits into four connected components, which are characterized by the sign before the monomial $y^{6}$ and the sign of the coefficient $\tau$. These sets are all $\mathfrak{G}$}-invariant, and the critical values of their functions at the $A_{2}$ points are constant along $\mathfrak{G}$}-orbits. It is easy to show that the polynomials with different combinations of these signs cannot be $\mathfrak{G}$}-equivalent to each other. The group ${\mathbb{Z}}_{2}^{2}$ of two involutions (4), (5) maps all these four sets into each other. The orbits of the group $\mathfrak{G}$} that contain the functions (8) with different values of the parameter $t$ are distinct because they have different critical values at the points of class $A_{2}$. If a polynomial of the form (1) belongs to one of these sets, then its entire orbit under the action of the group (6) also belongs to it. $\Box$
Lemma 18.
The subset of the space $\Phi_{3}$ consisting of polynomials having a singularity of class $D_{8}^{-}$ with zero critical value, and also a $A_{2}$ singularity, is the union of four connected components diffeomorphic to ${\mathbb{R}}^{7}$ and mapped to each other by the involutions ((4)) and ((5)). Two of these components consist of polynomials with the negative critical values at the $A_{2}$ points, and the other two consist of polynomials with the positive critical values. Each of these four components is swept out by a one-parametric family of orbits of the group $\mathfrak{G}$}. The intersection of each component with the parameter space of deformation ((2)) is an orbit of the group $\mathbb{R}$$_{+}$ of positive quasihomogeneous dilations ((6)).
Proof. Each function $f$ of class $\Phi_{3}$ with a critical point of class $D_{8}^{-}$ and critical value 0 can be reduced by the action of the group $\mathfrak{G}$} to the form
(9) $x^{3}+bx^{2}y^{2}\pm xy^{4}+\alpha x^{2}y.$
Indeed, the 2-jet of $f$ at this point is trivial, and the 3-jet has the form $q(x-\lambda y)^{2}(x-\nu y)$ for some $q>0$ and $\lambda\neq\nu$. The substitution $\tilde{x}=x-\lambda y$ and a dilation reduces this 3-jet to the form $\tilde{x}^{3}+\alpha\tilde{x}^{2}y$, $\alpha\neq 0$. The resulting function has zero coefficient at the monomial $y^{4}$: otherwise it would be of class $D_{5}$. The lower quasihomogeneous part with weights $\deg x=2,\deg y=1$ of the obtained polynomial should have the form $y(x-\theta y^{2})^{2}$, because otherwise the function is of the class $D_{6}$. So, this quasihomogeneous part can be reduced to the monomial $x^{2}y$ by a diffeomorphism of class $\mathfrak{G}$} (keeping the entire function in the class $\Phi_{3}$). The coefficient of the resulting polynomial at the monomial $y^{6}$ should vanish: otherwise the critical point is of class $D_{7}$. Finally, its coefficient at the monomial $xy^{4}$ is not zero, because otherwise the Milnor number would be greater than $8$. Dilating the coordinate $y$, we get the form (9).
The sum of Milnor numbers of all critical points of the obtained function is equal to ten, so there are only two critical points outside the origin (counting with the multiplicities). The system of equations $f^{\prime}_{x}=0=f^{\prime}_{y}$ for these two points implies easily the condition
$\pm 48y^{2}=12b^{2}y^{2}+20b\alpha y+7\alpha^{2}$
on their $y$-coordinate. If these two critical points coincide (at a point of class $A_{2}$), then $b^{2}\pm 21=0$, so the sign $\pm$ in (9) is equal to - and $b=\sqrt{21}$. The concluding arguments are the same as in the proof of Lemma 17. $\Box$
Corollary 19.
The space of polynomials ((1)) contains exactly two polynomials having a critical point of class $E_{8}$ with zero critical value and a critical point of class $A_{2}$ with critical value 1. The space of polynomials ((2)) contains exactly two polynomials having a critical point of class $D_{8}^{-}$ with zero critical value and a critical point of class $A_{2}$ with critical value 1. In both cases, these two points are mapped to each other by the involution ((4)).
Proposition 9 follows from this corollary in the same way as Proposition 3 of [15] was deduced from Lemmas 4 and 5 and Proposition 22. Namely, using the Lyashko–Looijenga covering, we can assume that our two Morse polynomials $f,\tilde{f}$ have equal sets of critical values. Using the diffeomorphisms of class $\mathfrak{G}$} we can also assume that they lie in the space of deformation (1) or (2). We draw a generic path in this space, connecting the polynomial $f$ with a point in the distinguished stratum $(E_{8},A_{2})$ or $(D_{8}^{-},A_{2})$ considered in Corollary 19. Using the Lyashko–Looijenga covering (and the coincidence of virtual Morse functions associated with $f$ and $\tilde{f}$), we can uniquely draw a path starting from the other polynomial $\tilde{f}$, repeating the sets of critical values of the functions of the first path and undergoing the same standard surgeries. The endpoint of the latter path can only either coincide with the endpoint of the first path, or be symmetric to it by the reflection (4). In a small neighborhood of these endpoints, the statement of Proposition 9 is trivial. By the continuity along our paths, this also holds for their starting points. $\Box$
5. Realization of some virtual Morse functions by polynomials
Almost all realizations given in this section are invariant under the involution (4) and hence prove the achirality of the corresponding components.
Remark 20.
In all our pictures, the $y$ axis is horizontal and $x$ axis is vertical and directed upwards.
5.1. $\Phi_{1}$
A virtual Morse function with ten real critical points, of which exactly one is a local maximum, can be realized by the polynomial
(10) $(x^{2}+y^{4}-8y^{2})(x+5(y+3/2)^{2}-5).$
A virtual Morse function with ten real critical points, of which exactly two are local maxima, can be realized by the polynomial
(11) $(x^{2}+y^{4}-2y^{2})(x+y^{2}-1),$
in particular the corresponding component is achiral. The zero sets of these two polynomials are outlined in the left parts of Figs. 3 and 4.
Virtual Morse functions with eight real critical points and exactly two (respectively, one, none) maximum points can be realized by polynomials whose zero level sets are shown in Fig. 18 left (respectively, center, and right).
Figure 18. Perturbations for $\Phi_{1}$ with eight real critical points
For the first of them, we can take the polynomial
(12) $x\left((y^{2}-1)^{2}+x^{2}-\varepsilon\right),\quad\varepsilon\in(0,1),$
For the second, we can take the polynomial
(13) $(x^{2}+y^{4}-1)(x+2y^{2}-A),\quad A\in(1,\sqrt{2}).$
To construct the third polynomial, we take the perturbation $x^{3}+y^{6}+\varepsilon(y^{5}+xy^{4})$, $\varepsilon>0$, of singularity $x^{3}+y^{6}$. It has one critical point of class $E_{8}$ and no other real critical points. Then we apply the standard perturbation of this $E_{8}$ singularity, as described on page 17 of [1].
Figure 19. Perturbations for $\Phi_{1}$ with six real critical points
The perturbation
(14) $x^{3}+y^{6}-\varepsilon xy^{2}$
of $J_{10}^{1}$ singularity $x^{3}+y^{6}$ has a critical point of class $D_{4}^{-}$, two local minima and no other critical points, see Fig. 19 (center). Additional perturbations of its $D_{4}^{-}$ critical point, shown in Fig. 19 left and right can be performed as indicated in [1], [14] preserving the symmetry with respect to the coordinate $y$. These perturbations realize virtual Morse functions with six real critical points, exactly one (respectively, none) of which are maxima. The perturbation
(15) $x^{3}+y^{6}-\varepsilon xy^{2}+\varepsilon^{3}x$
of the same function (14) keeps both of its minima and splits the $D_{4}^{-}$ point into two saddlepoints and two imaginary critical points, see Fig. 20 (center).
The perturbation
(16) $x(x^{2}+y^{4}+\varepsilon y^{2}-\varepsilon^{3})$
of singularity $x^{3}+xy^{4}$ of class $J_{10}^{1}$ has four real critical points: one maximum, one minimum and two saddles, see Fig. 20 (left).
The perturbation
(17) $x^{3}+y^{6}+\varepsilon(y^{2}-x)$
has only two real critical points: a minimum and a saddle.
Figure 20. Perturbations for $\Phi_{1}$ with four or two critical points
Finally, the perturbation
(18) $x^{3}+y^{6}+\varepsilon x$
has no real critical points.
5.2. $\Phi_{3}$
Two polynomials of class $\Phi_{3}$ with ten real critical points and one maximum point, realizing the D-graphs shown in Figs. 10 and 11, can be chosen as
(19) $x(x+3y^{2}-2)(x+y^{2}-1)$
and
(20) $x(x+3y(y+1)-1)(x+\frac{1}{3}(y-1)(y-3)-1).$
Their zero sets are shown in the left parts of the same figures; the first of them is invariant under the involution (4). The polynomial of class $\Phi_{3}$ with ten real critical points, exactly two of which are the maxima, can be chosen as follows:
(21) x(x+2(y+1)(y2)1/2)(x+(y1)(y+2)1),x(x+2(y+1)(y-2)-1/2)(x+(y-1)(y+2)-1),
see the left part of Fig. 12.
Figure 21. $\Phi_{3}$, eight critical points, one maximum, 66906
A polynomial of class $\Phi_{3}$ with eight real critical points, exactly one of which is a maximum, is given by
(22) $x(x+(y-1)^{2}-2)(x+(y+1)^{2}-2)+\varepsilon y^{6}\ .$
The zero level set of this polynomial without the term $\varepsilon y^{6}$ is shown in Fig. 21.
Figure 22. $\Phi_{3}$, eight critical points, no maxima
Two polynomials representing two different isotopy classes with eight real critical points, none of which are the maxima, are shown in Fig. 22. The first of these polynomials is
(23) $x(x+(y-2)^{2}-2)(x+(y+2)^{2}-2)+\varepsilon y^{6},$
it realizes the class with Card = 131148. The left picture of Fig. 22 shows the zero set of this polynomial without the term $\varepsilon y^{6}$. The second virtual component has Card = 82350 and probably has no symmetric realizations. Nevertheless it is achiral, see § 6.1.The zero set of some polynomial representing it is given in Fig. 22 (right). A realization of this picture is as follows: first we take the perturbation
$\frac{1}{3}(x^{3}-xy^{4})-\varepsilon x^{2}y$
of the original $J_{10}^{3}$ singularity $\frac{1}{3}(x^{3}-xy^{4})$, and then perturb its single real critical point of type $D_{8}^{-}$ as shown on p. 15 of [1].
An isotopy invariant proving that these two Morse functions indeed belong to different isotopy classes is as follows. For any function $f$ of class $\Phi_{3}$ with exactly three local minima, a unique parabola (or line) of the form $x=ay^{2}+by+c$ exists that passes through these three points. The restriction of $f$ to this curve is a polynomial of degree at most six in the coordinate $y$. Since it has three minima, it must be of degree exactly six. Therefore, the coefficient $a$ is not a root of the polynomial $\alpha t^{3}+\beta t^{2}+\gamma t+\delta$ where $\alpha x^{3}+\beta x^{2}y^{2}+\gamma xy^{4}+\delta y^{6}$ is the principal quasihomogeneous part of $f$. Thus, the position of $a$ among three roots of this polynomial is an invariant of the isotopy class. For the left picture of Fig. 22 the coefficient $a$ does not separate these roots, while for the right picture it separates one root from the other two.
Figure 23. $\Phi_{3}$, six critical points; one maximum (left) and no maxima (right)
The isotopy components of class $\Phi_{3}$ with exactly six real critical points, exactly one or none of which is a local maximum, are represented by the polynomials
(24) $x(x-y^{2}+4)(x+y^{2}+1)$
or
(25) $(x+\varepsilon)\left(2x^{2}+5xy^{2}+2y^{4}-8\frac{\sqrt{2}}{3}x-10\frac{\sqrt{2}}{3}y^{2}+\frac{25}{9}\right)$
with sufficiently small $\varepsilon>0$. Their zero-level sets are shown in Fig. 23. The degeneration of the second of these polynomials at $\varepsilon=0$ has only two real critical points of type $A_{3}$.
Figure 24. $\Phi_{3}$. Four critical points, Card=21410 (left); two critical points, Card=14778 (right)
A polynomial of class $\Phi_{3}$ with exactly four real critical points and no local maxima is given by
(26) $x(x-y^{2}+1)(x+y^{2}+3),$
see Fig. 24 (left). A polynomial with only two real critical points is given by
(27) $x(x-y^{2}-1)(x+y^{2}+1),$
see Fig. 24 (right).
6. Chirality of virtual Morse functions. Proof of Theorem 15
6.1. Chirality cocycle and proof of achirality statements of Theorem 15.
Consider a virtual component on the formal graph of class $\Phi_{1}$ or $\Phi_{3}$.
Denote by $\mathfrak{H}$} the one-dimensional simplicial cochain of this virtual component with coefficients in ${\mathbb{Z}}_{2}$, which takes the non-zero value on the edges corresponding to the elementary operation $s2$ of changing the order of two neighboring real critical values, both of which are achieved at minimum points.
Theorem 21.
If all virtual Morse functions of a virtual component have exactly two or three minima, then this component is chiral if and only if the cocycle $\mathfrak{H}$} is trivial on it.
Proof. Let $N=2$ or $3$ be this number of minima of all these functions. Consider a generic path in the space of Morse polynomials of class $\Phi_{1}$ or $\Phi_{3}$ that connects two mutually symmetric polynomials $f(x,y)$ and f(x,y)f(x,-y). The endpoints of this path are associated with the same virtual Morse function; therefore, the associated path in the formal graph is a cycle. Polynomials of class $\Phi_{1}$ or $\Phi_{3}$ cannot have two minima on the same vertical line. Therefore, our path preserves the order of the projections of the minimum points to the $y$ axis. On the other hand, for any $i\in[1,N]$ the critical value of the $i$th in this order minimum point of the starting polynomial of this path is equal to the critical value of the (Ni)(N-i)-th minimum point of the final polynomial of the path. Thus, our cycle contains an odd number of $s2$ surgeries that change the order of the critical values at the minimum points.
Conversely, given a cycle in a virtual component of the formal graph, let us realize it by a path in the space $\Phi_{1}$ or $\Phi_{3}$ of real polynomials starting from an arbitrary vertex of the cycle. That is, the path should start with a polynomial associated with this vertex and consecutively experience all surgeries encoded by the edges of the cycle. The existence of such a path is ensured by Proposition 7. If the cocycle $\mathfrak{H}$} takes the non-zero value on the initial cycle, then the isotopy classes of the endpoints of this realized path are different. $\Box$
Remark 22.
We could formulate the same criterion in terms of the critical values at the maximum points. Therefore, the other cocycle defined in this virtual component as $\mathfrak{H}$} but in the terms of the maxima instead of the minima is homologous to $\mathfrak{H}$}.
Almost all virtual components for which Theorem 15 claims their achirality are represented by polynomials that are symmetric with respect to the involution (4), see (11), (12), (13), (14), (15), (16), (17), (18), (19), (22), (23), (24), (25), (26), and (27). The remaining two virtual Morse functions are of class $\Phi_{3}$ with eight real critical points and Card = 82350; one of them has no maxima (see Fig. 22 right) and the other is formed from it by the virtual analog of involution (5). The simplicial cocycle $\mathfrak{H}$} takes the non-zero value on the cycle $A\to B\to D\to C\to A$ in the corresponding virtual component, see Fig. 2 and Example 5. Indeed, this cocycle takes the non-zero value only on the edge $[A,C]$ of this cycle. This proves the achirality of these two virtual components as well.
Remark 23.
This cycle was found by a small suspension of our program.
6.2. Proof of chirality statements of Theorem 15
Proposition 24.
If a virtual component of class $\Phi_{1}$ or $\Phi_{3}$ consists of virtual functions with ten real critical points and is achiral, then its $D$-graph has an automorphism, whose restriction to the set of vertices corresponding to minimum points ((or to the set of vertices corresponding to maximum points)) is an involution with at most one fixed point.
Proof. Let $f(x,y)$ be a generic Morse polynomial in an achiral isotopy class with ten real critical points. There are two one-to-one correspondences between its critical points and the critical points of the polynomial f(x,y)f(x,-y). One correspondence preserves the critical values of the critical points, and the other is obtained by continuing over the path connecting the two polynomials in the isotopy class. The composition of these one-to-one correspondences of critical points (and hence of the corresponding vertices of the associated D-graphs) extends to an automorphism of the D-graph of $f$. This automorphism maps the vertex corresponding to the minimum point with the $i$th value from above of coordinate $y$ to the vertex corresponding to the point with the $i$th value from below. Similarly, it permutes the points of maxima. In particular, it defines an involutive permutation of each of these two sets of points with at most one fixed point in each. $\Box$
This proposition implies the chirality of all virtual components with ten real critical points, except for those whose achirality follows from formulas (11) and (19).
It remains to prove the chirality of the component of class $\Phi_{1}$ with eight real critical points, none of which are maxima, see Fig. 18 (right). (The criterion of Theorem 21 does not work for it, since the number of minima is not 2 or 3).
Definition 13.
A level section on a virtual component of the formal graph is a natural-valued function on the set of its vertices, such that
a) for any edge not of type $s2$ this function takes equal values at its endpoints,
b) for any edge corresponding to the $s2$ surgery at which the $i$th and $(i+1)$th real critical values meet and overtake each other, the values at its endpoints are either equal to each other and not equal to $i$ or $i+1$, or one of these values is equal to $i$ and the other to $i+1$.
Theorem 25.
Suppose that $\varphi$ is an arbitrary vertex of an achiral virtual component with an even positive number of minima and the $i$th ((in the ascending order of critical values)) critical point of the virtual function $\varphi$ is a minimum. Then there does not exist a level section of this component, whose value at $\varphi$ is $i$.
Proof. Suppose that such a level section exists. Let $f(x,y)$ be a generic Morse polynomial associated with $\varphi$. Consider again a generic path in the space of Morse functions of class $\Phi_{1}$ or $\Phi_{3}$ connecting the function $f$ with its mirror image f(x,y)f(x,-y), and the path in our virtual component starting from $\varphi$ and consisting of edges corresponding to virtual surgeries describing all real surgeries along the previous path. The latter path is then a cycle. The continuation of the level section’s value along this cycle leads to a number different from the number at its starting point: a contradiction. $\Box$
On the other hand, such a level section for the investigated virtual component was constructed by our program. This finishes the proof of Theorem 15. $\Box$
7. Proof of Theorem 2
Recall that the canonical Dynkin graphs of the real singularities $D_{4}^{+}$, $D_{6}^{+}$, and $D_{8}^{+}$ are given in Fig. 25.
Figure 25. Real Dynkin graph $D_{8}^{+}$
The canonical Dynkin graphs of the real singularity $D_{2k}^{-}$ is just the standard $D_{2k}$ graph. The canonical Dynkin graphs of the other real simple singularities, $A_{k}$, $D_{2k+1}$, $E_{6}$, $E_{7}$, and $E_{8}$, are also the same as in the classical “complex” theory.
All singularities that appear as perturbations of a $J_{10}$ singularity are simple (or equal to $J_{10}$).
Proposition 26.
[…]
a)) all edges connecting vertices of different subsets are directed from vertices of one subset to vertices of the other, and
b)) the edges of our D-graph, both of whose vertices belong to one of these subsets, form the canonical Dynkin graph of the corresponding real simple singularity $\Xi$ or $\tilde{\Xi}$.
Proof. A proof of the “only if” part repeats the proof of Proposition 26 in [15]. The “if” part essentially follows from the same considerations with the Lyashko–Looijenga covering. (It was not formulated in [15] only because the realizations of all necessary splittings were explicitly presented there). $\Box$
Theorem 2 is a direct corollary of this proposition and the lists of D-graphs given in Theorems 13 and 14. $\Box$